OpenStudy (anonymous):

What is the solution of the equation? 6x – 3 = 5x – 5

OpenStudy (anonymous):

Bring x's to one side, and everything else to the other side 6x - 5x = -5 + 3 x = -2

OpenStudy (anonymous):

6(-2)-5(-2)+3

OpenStudy (anonymous):

put that in the calculatoe and u have your answer